OpenStudy (anonymous):

find the slope and y-intercept of the line 4x+2y=12

OpenStudy (mathhelp346):

turn it into the y=mx+b format

OpenStudy (mathhelp346):

subtract 4x and get 2y=-4x+12

OpenStudy (anonymous):

thats the answer

OpenStudy (mathhelp346):

yes. -4 is the slope and 12 is the y intercept

OpenStudy (anonymous):

\[-2x+2<6\]

OpenStudy (mathhelp346):

turn the sign into an equal sign and graph

OpenStudy (anonymous):

how

OpenStudy (mathhelp346):

-2x+2=6

OpenStudy (mathhelp346):

like i did with the last problem, turn it into the y=mx+b format

OpenStudy (mathhelp346):

then you'll get your slope and y-intercept

OpenStudy (anonymous):

-2(x+3)+6x